OpenStudy (agent_a):

A particular brand of gasoline has a density of 0.737 g/mL at 25°C. How many grams of this gasoline would fill a 13.4 gal tank?

OpenStudy (agent_a):

This is what I did but it was wrong: \[Density = \frac{ mass }{ volume }\] \[13.4 gal * 3.79 = 50.786 \frac{ L }{ gal }\] 50.786 \[(50.786 \frac{ L }{ gal })(\frac{ 10^{-3} mL}{ 1L }) = 0.050786 \frac{ mL}{ gal }\] \[(0.737 \frac{ g }{ mL })(0.050786 \frac{ mL }{ gal }) = 0.037429282 \frac{ g }{ gal }\] How do I take the "gal" out of the units of my final answer?